# thanks: A footnote to the article title

### v2

@@ Abstract
As the workhorse of electronic structure methods, density functional theory (DFT) has become essential for molecule and materials discovery, deploying databases covering vast chemical and composition spaces, predicting chemical and electrochemical reactions, and developing machine learning potentials. The extensive application of DFT in chemistry and materials science requires predictions achieving ”chemical accuracy,” a goal hindered by the unknown functional form for the exact exchange and correlation (XC) energy. A recent meta-generalized gradient approximation, the restored-regularized strongly constrained and appropriately normed (r2SCAN) XC functional, fulfils 17 exact constraints of the XC energy, and has significantly boosted the accuracy of predictions for molecules and materials. However, r2SCAN still appears inadequate at predicting material properties of open d and f transition-metal strongly correlated compounds, including band gaps, magnetic moments, and oxidation energies. Prediction inaccuracies of r2SCAN arise from functional and density-driven errors, mainly resulting from the DFT self-interaction error. Here, we propose a novel method termed r2SCANY@r2SCANX to mitigate the pernicious self-interaction error of XC functionals for the accurate simulations of electronic, magnetic, and thermochemical properties of transition metal oxides. r2SCANY@r2SCANX utilizes different fractions of exact Hartree-Fock exchange: X to define the electronic density, and Y to set the density functional approximation for the energy. We show that r2SCANY@r2SCANX simultaneously addresses functional-driven and density-driven inaccuracies, mitigating the self-interaction error in DFT. Building just on 1 (or maximum 2) universal parameters, we demonstrate that r2SCANY@r2SCANX improves upon the r2SCAN predictions for 18 highly correlated oxides and even outperforms the highly parameterized DFT(r2SCAN)+U method –the state-of-the-art approach to predict strongly correlated materials. We demonstrate that the typical O2 overbinding inaccuracies of the local density and the gradient XC approximations, with errors ranging between –2.2 and –1.0 eV/O2, reduce to –0.3 eV/O2 for r2SCAN, and further reduce to less than $\sim$0.03 eV/O2 using any combination of X in r2SCAN10@r2SCANX, including X=0. Prediction uncertainties for oxidation energies and magnetic moments of transition metal oxides are significantly reduced by r2SCAN10@r2SCAN50 and band gaps with r2SCAN10@r2SCAN. r2SCAN10@r2SCAN50 diminishes the density-driven error present in r2SCAN and r2SCAN10. We demonstrate that the computationally efficient r2SCAN10@r2SCAN is nearly as accurate as the global hybrid r2SCAN10 for oxidation energies. This suggests that accurate energy differences can be achieved through rate-limiting self-consistent iterations and geometry optimizations using the efficient r2SCAN. Subsequently, a more expensive nonlocal functional, such as a hybrid or self-interaction correction, can be applied in a fast-to-execute single post-self-consistent calculation, as in r2SCAN10@r2SCAN. Compared to computationally demanding r2SCAN10, r2SCAN10@r2SCAN appears 10 to 300 times faster.
@@ I Introduction
In this context, the strongly constrained and appropriately normed (SCAN)19 and its regularized and computationally efficient version, r2SCAN,20 are highly accurate meta-GGA functionals. Both SCAN and r2SCAN satisfy 17 known exact constraints in constructing XC functionals, ensuring a well-balanced description of XC effects for a wide range of systems.21 Another constraint-based meta-GGA is LAK, which accurately describes the electronic bonding (at least for main group semiconductors) and band gaps of molecules and solids.22 SCAN and r2SCAN have done admirably well in improving the quality of predictions of molecules and materials relative to LSDA and GGA.23, 24, 25, 26, 27, 28 Note that most databases discussed above rely entirely on the predictive capabilities of GGA, and ad hoc corrections, including GGA+U,29, 30 and more recently r2SCAN.31
@@
Pragmatic but material- and property-dependent solutions to address the SIE in GGA and meta-GGA functionals involve parametrizing ad hoc on-site Hubbard U corrections, such as GGA(LSDA)+U,37 and r2SCAN(SCAN)+U.25, 26, 27, 29, 30, 38, 39, 40 Although numerically accurate and computationally affordable, these XC functionals+U approaches generally lack transferability as the U parameters depend on material chemistry, dimensionality, and the oxidation states of transition metals. Numerous strategies exist to fit the U values to the thermochemical data of redox reactions,25, 26, 27, 29, 30, 39, 40 band gaps.37, 41 U values can also be derived from linear response,42, 43 or even with machine learning models.44, 45 A persistent issue exists: when a U-specific value is applied to oxidation energies, it is not evident that this same numerical value of U will accurately replicate the band gaps (or magnetic moments) under similar conditions in materials.29, 30
@@ II Results
II.1 Improving Ab initio Predictions of meta-GGA-type r2SCAN Functionals
All XC functionals inherently suffer from functional-driven and density-driven errors (Eq. 1). SIE, a significant part of both errors, leads to inaccurate predictions of energetics in molecules and materials. LSDA, GGA, meta-GGA XC functionals struggle to accurately describe properties, such as reaction energies, inter-atomic charge transfer, and electronic structure, particularly in systems with strongly localized d (or f) electrons, such as MiOjs.10, 25, 26, 29, 30, 37, 38, 39, 39, 94, 95, 96, 97, 98
To address the SIE, we present the generalized r2SCANY@r2SCANX approach, which builds upon r2SCAN. In r2SCANY@r2SCANX, the SIE is addressed by independently tuning the fraction of exact HF exchange directly in the r2SCAN functional definition used to evaluate the total energy of a given set of orbitals and the r2SCAN functional used to compute these orbitals. Specifically, in r2SCANY@r2SCANX, Y is the percent of HF exact exchange appearing in the XC potential of r2SCAN to correct the functional-driven error. In contrast, in r2SCANY@r2SCANX, X is the percent of the exact Hartree-Fock exchange used in the r2SCAN functional to determine the electronic orbitals and potentially correct for the density-driven error. By independently varying the Xand Yparameters in r2SCANY@r2SCANX, the proposed method allows for a more systematic correction of the functional- and density-driven errors inherent to r2SCAN’s formulation, with the aim of improving the accuracy in describing transition metal oxides and potentially other correlated systems.
@@ II.1 Improving Ab initio Predictions of meta-GGA-type r2SCAN Functionals
where, in Eq. 2 $n^{X}$ is the self-consistent electron density obtained from the r2SCANX Hamiltonian, which includes Xexact exchange. Therefore, the r2SCANY@r2SCANX definition incorporates different fractions of exact exchange in both the functional definition and the charge density, ensuring a more systematic correction of SIE. Note, that replacing a small fraction of r2SCAN exchange by the same fraction of exact Hartree-Fock exchange should have minimal or no effect on the satisfaction of exact DFT constraints of this meta-GGA.
@@
Figure 1: Schematically connecting the self-consistent and non-self-consistent approaches required in the r2SCANY@r2SCANX methods. Self-consistent functionals used in this work are LSDA, PBE, r2SCAN, r2SCAN10, and LAK. Non-self-consistent hybrid functionals, including a fraction of exact HF exchange, such as r2SCAN@HF, r2SCAN10@r2SCAN, r2SCAN@r2SCAN50, r2SCAN10@r2SCAN50, r2SCAN@LAK, r2SCAN10@LAK, require a self-consistent step to generate orbitals that are evaluated non-self-consistently.
@@
II.2 r2SCAN Predictions of Transition-metal Oxides Properties
@@
Table 1 summarizes the experimentally predicted lattice parameters, magnetic moments, and band gaps of all materials using r2SCAN and r2SCAN+U (with U parameters calibrated in Ref. 39). All calculations reported in Table 1 used r2SCAN geometry-optimized structures, including the relaxation of atomic positions, volumes, and lattice shapes. Note, for the r2SCAN+U data set, the same functional was used for geometry relaxation (atomic positions, volumes, and shapes).
Table 1: r2SCAN and r2SCAN+U predicted lattice parameters, on-site magnetic moments (in $\mu$B), and band gaps (in eV) of MiOjs. Experimental quantities are marked as Exp. All calculations employ the experimentally-determined magnetic orderings (M.O.) and referenced in the magnetic moment (M.M.) column. Antiferro magnetic (AFM), ferrimagnetic (FEM), ferromagnetic (FM), and non-magnetic orderings (NM) are indicated.
@@
II.3 Benchmarking Oxidation Enthalpies of MiOjs with r2SCANY@r2SCANX Functionals
@@
where, $E\left(\mathrm{MO_{z}}\right)^{\mathrm{r^{2}SCANY@r2SCANX}}$, $E\left(\mathrm{MO_{x}}\right)^{\mathrm{r^{2}SCANY@r2SCANX}}$, $E\left(\mathrm{O_{2}}\right)^{\mathrm{r^{2}SCANY@r2SCANX}}$, are the r2SCANY@r2SCANX total energies of the oxidized, reduced MiOjs, and $\mathrm{O_{2}}$ gas, respectively. Our DFT predictions are compared to the experimental oxidation enthalpies (at 298 K, and 1 atm) of $\mathrm{MO_{x}}$ and $\mathrm{MO_{z}}$ (Supplementary Table 1) extracted from experimental reports.149, 150, 151 Thermochemical table provide oxidation energies earnest estimates of the corresponding 95% confidence intervals of MiOjs. In addition, here we approximated the oxidation enthalpies by the DFT total energies by imposing $\Delta H_{0}\approx\Delta E$, thus neglecting the pV contributions, which are expected to be minimal.
@@
In Fig. 2c, r2SCAN exhibits a mean absolute error (MAE) of $\sim$1.19 eV/O2 and a root mean squared error (RMSE) of $\sim$1.42 eV/O2 in predicting oxidation enthalpies in our dataset. r2SCAN systematically makes oxidation energies too negative across all tested cases. This is largely a functional-driven error. The r2SCAN+U, which implements an ad hoc U correction.29, 39, 152 The U parameters used here are from Ref. 39. When applied to the 15 oxidation reactions, r2SCAN+U yields significantly milder deviation from experimental data, with an MAE of $\sim$0.57 eV/O2 and an RMSE of $\sim$0.66 eV/O2.
@@
When applying r2SCAN@HF and utilizing the HF electronic density for the r2SCAN total energy evaluation, the resultant MAE for our dataset is 1.69 eV/O2 and an RMSE of 2.03 eV/O2. This indicates a decline of approximately 40% in accuracy compared to standard r2SCAN. Similarly to r2SCAN@HF, one can reduce the fraction of HF exact exchange to 50% as in r2SCAN@r2SCAN50 while producing a potentially lower density-driven error than r2SCAN.70, 153 This method gives our dataset an MAE of 0.86 eV/O2 (Fig. 2c) and an RMSE of 0.96 eV/O2.
Building upon r2SCAN@HF, a novel r2SCAN derivative proposed here is the r2SCAN10@r2SCAN, which iterates to self-consistency with the less expensive r2SCAN, requires only a single total-energy evaluation (not requiring a complete self-consistent field electronic relaxation), with the more costly global hybrid r2SCAN10. Supposedly, r2SCAN10@r2SCAN corrects functional-driven errors by introducing 10% exact exchange in the functional while using r2SCAN orbitals.
@@
The XC functional LAK is expected to predict more accurate band gaps, which are closely related to charge transfer processes in materials. For this reason, LAK is generally believed to yield improved electron densities compared to r2SCAN,22, 154 and hence, speculatively, r2SCAN@LAK and r2SCAN10@LAK should predict oxidation energies in better agreement with experimental data. However, for our dataset, the performance of r2SCAN@LAK (MAE: $\sim$1.21 eV/O2) appears comparable to that of r2SCAN (MAE: 1.19 eV/O2), and similarly, r2SCAN10@LAK (MAE: 0.81 eV/O2) shows no evident improvements over r2SCAN10@r2SCAN (MAE: 0.74 eV/ O2).
@@
II.4 Achieving Optimal Fractions of Exact Exchange in the r2SCANY@r2SCANX Formulations

### v3

# Reducing Self-Interaction Error in Transition-Metal Oxides with Different Exact-Exchange Fractions for Energy and Density

@@ Abstract
As the workhorse of electronic structure methods, density functional theory (DFT) has become essential for molecule and materials discovery, deploying databases covering vast chemical and composition spaces, predicting chemical and electrochemical reactions, and developing machine learning potentials. The extensive application of DFT in chemistry and materials science requires predictions achieving ”chemical accuracy,” a goal hindered by the unknown functional form for the exact exchange and correlation (XC) energy. A recent meta-generalized gradient approximation, the restored-regularized strongly constrained and appropriately normed (r2SCAN) XC functional, fulfils 17 exact constraints of the XC energy, and has significantly boosted the accuracy of predictions for molecules and materials. However, r2SCAN still appears inadequate at predicting material properties of open d and f transition-metal strongly correlated compounds, including band gaps, magnetic moments, and oxidation energies. Prediction inaccuracies of r2SCAN arise from functional and density-driven errors, mainly resulting from the DFT self-interaction error. Here, we propose a novel method termed r2SCANY@r2SCANX to mitigate the pernicious self-interaction error of XC functionals for the accurate simulations of electronic, magnetic, and thermochemical properties of transition metal oxides. r2SCANY@r2SCANX utilizes different fractions of exact Hartree-Fock exchange: X to define the electronic density, and Y to set the density functional approximation for the energy. We show that r2SCANY@r2SCANX simultaneously addresses functional-driven and density-driven inaccuracies, mitigating the self-interaction error in DFT. Building just on 1 (or maximum 2) universal parameters, we demonstrate that r2SCANY@r2SCANX improves upon the r2SCAN predictions for 18 highly correlated oxides and even outperforms the highly parameterized DFT(r2SCAN)+U method –the state-of-the-art approach to predict strongly correlated materials. We demonstrate that the typical O2 overbinding inaccuracies of the local density and the gradient XC approximations, with errors ranging between –2.2 and –1.0 eV/O2, reduce to –0.3 eV/O2 for r2SCAN, and further reduce to less than $\sim$0.03 eV/O2 using any combination of X in r2SCAN10@r2SCANX, including X=0. Prediction uncertainties for oxidation energies and magnetic moments of transition metal oxides are significantly reduced by r2SCAN10@r2SCAN50 and band gaps with r2SCAN10@r2SCAN. r2SCAN10@r2SCAN50 diminishes the density-driven error present in r2SCAN and r2SCAN10. We demonstrate that the computationally efficient r2SCAN10@r2SCAN is nearly as accurate as the global hybrid r2SCAN10 for oxidation energies. This suggests that accurate energy differences can be achieved through rate-limiting self-consistent iterations and geometry optimizations using the efficient r2SCAN. Subsequently, a more expensive nonlocal functional, such as a hybrid or self-interaction correction, can be applied in a fast-to-execute single post-self-consistent calculation, as in r2SCAN10@r2SCAN. Compared to computationally demanding r2SCAN10, r2SCAN10@r2SCAN appears to be 12 to 165 times faster for the iron oxides in our test set.
@@ I Introduction
In this context, the strongly constrained and appropriately normed (SCAN)19 and its regularized and computationally efficient version, r2SCAN,20 are highly accurate meta-GGA functionals. Both SCAN and r2SCAN satisfy 17 known exact constraints in constructing XC functionals, ensuring a well-balanced description of XC effects for a wide range of systems.21 Another constraint-based meta-GGA is LAK, which accurately describes the electronic bonding (at least for the main group molecules) and band gaps (for main group semiconductors).22 SCAN and r2SCAN have done admirably well in improving the quality of predictions of molecules and materials relative to LSDA and GGA.23, 24, 25, 26, 27, 28 Note that most databases discussed above rely entirely on the predictive capabilities of GGA, and ad hoc corrections, including GGA+U,29, 30 and more recently r2SCAN.31
@@
Pragmatic but material- and property-dependent solutions to address the SIE in GGA and meta-GGA functionals involve parametrizing ad hoc on-site Hubbard U corrections, such as GGA(LSDA)+U,37 and r2SCAN(SCAN)+U.25, 26, 27, 29, 30, 38, 39, 40 Although numerically accurate and computationally affordable, these XC functionals+U approaches generally lack transferability as the U parameters depend on material chemistry, dimensionality, and the oxidation states of transition metals. Numerous strategies exist to fit the U values to the thermochemical data of redox reactions,25, 26, 27, 29, 30, 39, 40 or band gaps.37, 41 U values can also be derived from linear response,42, 43 or even with machine learning models.44, 45 A persistent issue exists: when a U-specific value is applied to oxidation energies, it is not evident that this same numerical value of U will accurately replicate the band gaps (or magnetic moments) under similar conditions in materials.29, 30
@@ II Results
II.A Improving Ab initio Predictions of meta-GGA-type r2SCAN Functionals
All XC functionals inherently suffer from functional-driven and density-driven errors (Eq. 1). SIE, a significant part of both errors, leads to inaccurate predictions of energetics in molecules and materials. LSDA, GGA, and meta-GGA XC functionals struggle to accurately describe properties, such as reaction energies, inter-atomic charge transfer, and electronic structure, particularly in systems with strongly localized d (or f) electrons, such as MiOjs.10, 25, 26, 29, 30, 37, 38, 39, 39, 94, 95, 96, 97, 98
To address the SIE, we present the generalized r2SCANY@r2SCANX approach, which builds upon r2SCAN. In r2SCANY@r2SCANX, the SIE is addressed by independently tuning the fraction of exact HF exchange directly in the r2SCAN functional definition used to evaluate the total energy of a given set of orbitals and the r2SCAN functional used to compute these orbitals. Specifically, in r2SCANY@r2SCANX, Y is the percent of HF exact exchange appearing in the XC energy of r2SCAN to correct the functional-driven error. In contrast, in r2SCANY@r2SCANX, X is the percentage of the exact Hartree-Fock exchange used in the r2SCAN functional to determine the electronic orbitals and potentially correct for the density-driven error. By independently varying the Xand Yparameters in r2SCANY@r2SCANX, the proposed method allows for a more systematic correction of the functional- and density-driven errors inherent to r2SCAN’s formulation, with the aim of improving the accuracy in describing transition metal oxides and potentially other correlated systems.
@@ II.A Improving Ab initio Predictions of meta-GGA-type r2SCAN Functionals
where, in Eq. 2 $n^{X}$ is the self-consistent electron density obtained from the r2SCANX Hamiltonian, which includes Xexact exchange. Therefore, the r2SCANY@r2SCANX definition incorporates different fractions of exact exchange in the energy and self-consistent electron density, ensuring a more systematic correction of SIE. Note, that replacing a small fraction of r2SCAN exchange by the same fraction of exact Hartree-Fock exchange should have minimal or no effect on the satisfaction of exact DFT constraints of this meta-GGA.
@@
Figure 1: Schematically connecting the self-consistent and non-self-consistent (non-SCF) approaches required in the r2SCANY@r2SCANX methods. Self-consistent functionals used in this work are LSDA, PBE, r2SCAN, r2SCAN10, and LAK. Non-self-consistent hybrid functionals, including a fraction of exact HF exchange, such as r2SCAN@HF, r2SCAN10@r2SCAN, r2SCAN@r2SCAN50, r2SCAN10@r2SCAN50, r2SCAN@LAK, r2SCAN10@LAK, require a self-consistent step to generate orbitals on which the energy is evaluated non-self-consistently.
@@
II.B r2SCAN Predictions of Transition-metal Oxides Properties
@@
Table 1 summarizes the experimentally predicted lattice parameters, magnetic moments, and band gaps of all materials using r2SCAN and r2SCAN+U (with U parameters calibrated in Ref. 39). All calculations reported in Table 1 used r2SCAN geometry-optimized structures, including the relaxation of atomic positions, volumes, and lattice shapes. Note, for the r2SCAN+U data set, the r2SCAN+U functional was used for geometry optimization (atomic positions, volumes, and shapes).
Table 1: r2SCAN and r2SCAN+U predicted lattice parameters, on-site magnetic moments (in $\mu$B), and band gaps (in eV) of MiOjs. Experimental quantities are marked as Exp. All calculations employ the experimentally-determined magnetic orderings (M.O.) and referenced in the magnetic moment (M.M.) column. Antiferro magnetic (AFM), ferrimagnetic (FEM), ferromagnetic (FM), and non-magnetic orderings (NM) are indicated. Spin-polarized calculations were employed for all FM, AFM, and FEM systems, whereas NM materials were treated with spin-restricted calculations.
@@
II.C Benchmarking Oxidation Enthalpies of MiOjs with r2SCANY@r2SCANX Functionals
@@
where, $E\left(\mathrm{MO_{z}}\right)^{\mathrm{r^{2}SCANY@r2SCANX}}$, $E\left(\mathrm{MO_{x}}\right)^{\mathrm{r^{2}SCANY@r2SCANX}}$, $E\left(\mathrm{O_{2}}\right)^{\mathrm{r^{2}SCANY@r2SCANX}}$, are the r2SCANY@r2SCANX total energies of the oxidized, reduced MiOjs, and $\mathrm{O_{2}}$ gas, respectively. Our DFT predictions are compared to negative experimental oxidation enthalpies (at 298 K, and 1 atm) of $\mathrm{MO_{x}}$ and $\mathrm{MO_{z}}$ (Supplementary Table 1) extracted from experimental reports.149, 150, 151 Thermochemical tables provide oxidation energies with an uncertainty chosen so that the true value should lie within the uncertainty range 95% of the time. Furthermore, the definition of the bounds of the mean absolute error used in theoretical work are narrower than the typical standard deviation set to a 95% level of confidence imposed in thermochemical tables. In addition, here we approximated the oxidation enthalpies by the DFT total energies by imposing $\Delta H_{0}\approx\Delta E$, thus neglecting the pV contributions, which are expected to be minimal.
@@
In Fig. 2c, r2SCAN exhibits a mean absolute error (MAE) of $\sim$1.19 eV/O2 and a root mean squared error (RMSE) of $\sim$1.42 eV/O2 in predicting oxidation enthalpies in our dataset. r2SCAN systematically makes oxidation energies too negative across all tested cases. This is largely a functional-driven error. The r2SCAN+U implements an ad hoc U correction.39 The U parameters used here are from Ref. 39. When applied to the 15 oxidation reactions, r2SCAN+U yields significantly milder deviation from experimental data, with an MAE of $\sim$0.57 eV/O2 and an RMSE of $\sim$0.66 eV/O2.
@@
When applying r2SCAN@HF and utilizing the HF electronic density for the r2SCAN total energy evaluation, the resultant MAE for our dataset is 1.69 eV/O2 and an RMSE of 2.03 eV/O2. This indicates a decline of approximately 40% in accuracy compared to standard r2SCAN. Similarly to r2SCAN@HF, one can reduce the fraction of HF exact exchange to 50% as in r2SCAN@r2SCAN50 while producing a potentially lower density-driven error than r2SCAN.70, 152 This method gives our dataset an MAE of 0.86 eV/O2 (Fig. 2c) and an RMSE of 0.96 eV/O2.
Building upon r2SCAN@HF, a novel r2SCAN derivative proposed here is the r2SCAN10@r2SCAN, which iterates to self-consistency with the less expensive r2SCAN, and requires only a single total-energy evaluation (not requiring a complete self-consistent field electronic relaxation) with the more costly global hybrid r2SCAN10. Supposedly, r2SCAN10@r2SCAN corrects functional-driven errors by introducing 10% exact exchange in the functional while using r2SCAN orbitals.
@@
The XC functional LAK is expected to predict more accurate band gaps, which are closely related to charge transfer processes in materials. For this reason, we expected that LAK might yield generally improved electron densities compared to r2SCAN,22, 153 and hence, speculatively, r2SCAN@LAK and r2SCAN10@LAK should predict oxidation energies in better agreement with experimental data. However, for our dataset, the performance of r2SCAN@LAK (MAE: $\sim$1.21 eV/O2) appears comparable to that of r2SCAN (MAE: 1.19 eV/O2), and similarly, r2SCAN10@LAK (MAE: 0.81 eV/O2) shows no evident improvements over r2SCAN10@r2SCAN (MAE: 0.74 eV/ O2).
@@
II.D Achieving Optimal Fractions of Exact Exchange in the r2SCANY@r2SCANX Formulations